# In-Context Freeze-Thaw Bayesian Optimization for Hyperparameter Optimization

### v1

@@ 1 Introduction
Hyperparameters are essential in deep learning (DL) to achieve strong model performance. However, due to the increasing complexity of these models, it is becoming more challenging to find promising hyperparameter settings, even with the help of hyperparameter optimization (HPO, Section 3.1) tools (Feurer & Hutter, 2019; Bischl et al., 2023). Traditional HPO techniques using Bayesian Optimization (BO) are unsuitable for modern DL because they treat the problem as a black box, making it computationally expensive requiring a full model training for each evaluation.
Recent research in HPO has shifted towards multi-fidelity methods (Li et al., 2017, 2020a; Falkner et al., 2018; Klein et al., 2020; Li et al., 2020b; Awad et al., 2021), utilizing lower fidelity proxies (e.g., training for fewer steps, using less data, smaller models) and only evaluating the most promising hyperparameter settings at the full fidelity. While these methods have potential, they often use coarse-grained fidelity spaces and rely on the rank correlation of performances across fidelities. Moreover, they do not fully utilize the anytime nature of algorithms such as checkpointing, continuation, and extrapolation. As a result, these methods struggle to allocate computational resources efficiently, leading to suboptimal performance in many scenarios.
The freeze-thaw BO method (FT-BO, Section 3.2), originally proposed by Swersky et al. (2014), is a promising approach to efficiently allocate computational resources by pausing and resuming the evaluation of different hyperparameter configurations. This method is an improvement over traditional grey-box methods as it dynamically manages resources, and recent implementations (Wistuba et al., 2022; Kadra et al., 2023) hold the state-of-the-art in the low-budget regime ($\sim$ 20 full function evaluations). However, these contemporary implementations have limitations. In particular, they rely on online learning to update the surrogate model at each step which can lead high computational overhead, instability, and complexity in managing additional hyper-hyperparameters. Moreover, they also suffer from strong assumptions about learning curves, which may not be applicable in all scenarios, resulting in overly confident incorrect predictions.
@@
In this work, we leverage prior-data fitted networks (Müller et al., 2022, PFNs) (Section 3.3), a Transformer-based meta-learning approach to Bayesian inference, to enhance freeze-thaw BO through in-context learning. Our PFN model (FT-PFN) infers the task-specific relationship between hyperparameter settings and their learning curves in a single forward pass, eliminating the need for online training during the search. Figure 1 compares learning curves extrapolation, including uncertainty, by our model (FT-PFN) and two baselines. Beyond demonstrating superior extrapolation quality, our model directly addresses key challenges in traditional HPO methods, lowering computational overhead and stabilizing the optimization process. The contributions of this paper are as follows.
  • We propose FT-PFN (Section 4.1), a new surrogate model for freeze-thaw BO, replacing online learning with in-context learning using PFNs. We train a single instance of FT-PFN model exclusively on synthetic data, generated from a curve prior (Section 4.1) designed to mimic realistic HPO learning curves.
@@
  • We combine FT-PFN with a novel acquisition function (MFPI-random, Section 4.2) and find that the resulting in-context freeze-thaw method (ifBO) yields a new state-of-the-art performance on the three benchmark suites for HPO for deep learning (LCBench, Taskset, PD1).
We will publicly release the code for the surrogate PFN training and reproducing experiments from this paper.
@@ 2 Related Work
uses low-cost approximations of the objective function, for example, by evaluating only a few epochs of model training. A notable approach in this category is Hyperband (Li et al., 2017), which iteratively allocates a similar budget to various candidate hyperparameter settings and retains only the most promising ones, to be evaluated at a higher budget. Hyperband was extended for efficient parallelization (Li et al., 2020a), Bayesian optimization (Falkner et al., 2018) and evolutionary search (Awad et al., 2021) and to incorporate user priors (Mallik et al., 2023). However, a key limitation of these multi-fidelity approaches is that they do not allow the continuation of previously discarded runs in the light of new evidence, resulting in a waste of computational resources. Additionally, their effectiveness heavily depends on a good manual choice of fidelity levels and a strong correlation between the ranks of hyperparameter settings at low and high fidelities (e.g., no crossings of learning curves).
@@
offers a promising solution to address the limitations of standard multi-fidelity and LCE-based HPO. Its ability to pause and resume optimization runs enables the efficient management of computational resources and ensures solid anytime performance. As a BO method, it incorporates a surrogate model to predict performance at higher fidelities and relies on the acquisition function to guide the search. The freeze-thaw concept was introduced by Swersky et al. (2014), who used an entropy search acquisition function and a Gaussian Process with exponential decay kernel for modeling learning curves. Wistuba et al. (2022) recently proposed DyHPO, an improved version that uses a learned deep kernel combined with a multi-fidelity-based Expected Improvement (EI) acquisition. Most recently, Kadra et al. (2023) introduced DPL, another surrogate model that utilizes a deep ensemble of power laws and Expected Improvement at the maximum budget as acquisition function. These Freeze-Thaw BO methods substantially improve upon standard multi-fidelity approaches, showing particularly strong performance in low-budget settings. Nevertheless, these methods share a common challenge: the necessity for online training of a surrogate model during the search, which can be computationally intensive and may cause optimization instabilities.
@@ 3.1 Hyperparameter Optimization (HPO)
Consider an iterative machine training pipeline with configurable hyperparameters $\lambda\in\Lambda$. E.g., when training a neural network using gradient descent we could configure the learning rate, weight decay, dropout rate, etc. Let $f(\lambda,t)$ be some measure of downstream performance (e.g., validation loss) of the model obtained, using hyperparameter settings $\lambda$, after $t$ iterations of training, that we would like to minimize. Further, assume that the training run for any $\lambda$ is limited to $b_{\text{max}}$ such iterations. HPO aims to find a hyperparameter setting producing the best model, i.e., $\lambda^{*}$: $(\lambda^{*},\cdot)\in\operatorname*{arg\!min}_{\lambda\in\Lambda,t\in\mathbb{N}_{0}}f(\lambda,t)$, within the limited total optimization budget of $B$ iterations. Note that in modern deep learning the budget available for HPO often does not allow us to execute more than a few full training runs. In this setting, the crux of HPO lies in allocating these limited training resources to the most promising hyperparameter settings, i.e., to find an allocation $\{b_{\lambda}\}_{\lambda\in\Lambda}$ with $\sum_{\lambda\in\Lambda}b_{\lambda}=B$ that minimizes $\min_{\lambda\in\Lambda,\,1\leq t\leq b_{\lambda}}f(\lambda,t)$.
3.2 Freeze-Thaw Bayesian Optimization (FT-BO)
@@
As discussed above, Swersky et al. (2014) proposed to address this challenge by following a fine-grained dynamic scheduling approach, allocating resources to configurations (and observing their performance) “one step at a time”. Here, one “step”, denoted as $\mathds{1}_{b}$, corresponds to one or more iterations of model training (e.g., one epoch). Algorithm 1 shows the freeze-thaw Bayesian optimization framework (FT-BO) which uses its history $H$, i.e., the various partial learning curves observed thus far in the current partial allocation, to fit a dynamic Bayesian surrogate model $\mathcal{M}{}$ that probabilistically extrapolates the partially seen performance of configuration $\lambda$ beyond $b_{\lambda}$ (Algorithm 1, line 4). Following the BO framework, it decides which of these to continue (or start if $b_{\lambda}=0$) using a dynamic acquisition policy $\mathcal{A}{}$ trading off exploration and exploitation (Algorithm 1, line 5). The crucial difference with traditional black box Bayesian Optimization lies in that resources in FT-BO are allocated one step, rather than one full training run, at the time. Note, FT-BO assumes training to be preemptive, i.e., that we can stop (freeze) a model training, and continue (thaw) it later; this assumption is reasonable in modern DL where checkpointing is common practice. Also note, that black box BO can be recovered as a special case, by setting $b_{\text{max}}$, the maximum number of steps allocated to any single configuration, to one, and $\mathds{1}_{b}$ to the maximum budget available for any single training run.
@@
   $f$: measure of model performance to be minimized,
@@
   $b_{\text{max}}$: maximal steps for any configuration $\lambda\in\Lambda$,
@@
Output: $\lambda\in\Lambda$, that obtained the best performance observed
Procedure: HPO($\Lambda$, $f$, $\mathds{1}_{b}$, $b_{\text{max}}$, $B$):
@@
2:
3:  while $|H|\cdot\mathds{1}_{b}<B$ do
4:   Train $\mathcal{M}{}$ on $H$ FT-PFN requires no model fitting
5:   $\lambda\xleftarrow{}$ $\mathcal{A}{}$($\Lambda,\mathcal{M}{},H,b_{\text{max}}$) select $\lambda$ to thaw
6:   $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n_{\lambda}\leftarrow{}\max\{z\}_{(\lambda,z,y)\in H}$$n_{\lambda}$: steps performed using $\lambda$
7:   $y\leftarrow{}f(\lambda,(n_{\lambda}+1)$$\cdot\mathds{1}_{b})$evaluate $f$ (thaw $\lambda$ for one step)
8:   $H\leftarrow H\cup\{(\lambda,n_{\lambda}+1,y)\}$
9:  end while
9:
10:  return $\lambda^{*}$: $(\lambda^{*},\cdot)\in\operatorname*{arg\!min}_{\lambda\in\Lambda,\,1\leq z\leq n_{\lambda}}f(\lambda,z\cdot\mathds{1}_{b})$
@@ 3.3 Prior-data Fitted Networks (PFNs)
As briefly discussed above, PFNs (Müller et al., 2022) are neural networks $q_{\theta}$ that are trained to do Bayesian prediction for supervised learning in a single forward pass. More specifically, let $D=D_{train}\cup\{(x_{test},y_{test})\}$ be a dataset used for training; the PFN’s parameters $\theta$ are optimized to take $D_{train}$ and $x_{test}$ as inputs and make predictions that approximate the posterior predictive distribution (PPD) of the output label $y_{test}$:
$q_{\theta}(x_{test},D_{train})\approx\mathbb{P}(y_{test}\,|\,x_{test},D_{train}),$
in expectation over datasets $D$ sampled from a prior $p(\mathcal{D})$ over datasets. At test time, the PFN does not update its parameters by gradient descent given a training dataset $D_{train}$, but rather takes $D_{train}$ as a contextual input, predicting the labels of unseen examples through in-context learning. The PFN is pretrained once for a specific prior $p(\mathcal{D})$ and used in downstream Bayesian prediction tasks without further fine-tuning. More specifically, it is trained to minimize the cross-entropy for predicting the hold-out example's label $y_{test}$, given $x_{test}$ and $D_{train}$:
$\begin{split}\ell_{\theta}=\mathbb{E}[\shortminus\text{log}\ q_{\theta}(y_{test}|x_{test},D_{train})]\\ \text{with \quad}\{(x_{test},y_{test})\}\cup D_{train}\sim p(\mathcal{D})\end{split}$ (1)
@@ 4 In-Context Freeze-Thaw BO (ifBO)
In this section, we describe ifBO, the in-context learned variant of the freeze-thaw framework that we propose as an alternative for the existing online learning implementations (Wistuba et al., 2022; Kadra et al., 2023). As can be seen in Algorithm 1 (line 4), the critical difference lies in the fact that we do not need to refit our surrogate model after every allocation step, but instead provide the full history of performance observations $H$ as input for prediction. By skipping the online refitting stage, we reduce computational overhead, code complexity, and hyper-hyperparameters. Note that within the FT-BO framework described in Section 3.2, our method is fully characterized by our choice of dynamic surrogate model (Section 4.1) and acquisition policy (Section 4.2).
@@ 4.1 Dynamic Surrogate Model (FT-PFN)
We propose FT-PFN a prior-data fitted network (Müller et al., 2022, PFNs) trained to be used as an in-context dynamic surrogate model in the freeze-thaw framework. As described in Section 3.3, PFNs represent a general meta-learned approach to Bayesian prediction, characterized by the data used for meta-learning. Following previous works using PFNs (Müller et al., 2022; Hollmann et al., 2023; Müller et al., 2023; Adriaensen et al., 2023; Dooley et al., 2023), we train the PFN only on synthetically generated data, allowing us to generate virtually unlimited data and giving us full control over any biases therein. We aim to train FT-PFN on artificial data ($D$) that resembles the real performance data we observe in the context of FT-BO, i.e., a collection of learning curves of training runs for the same task, but using different hyperparameter settings, i.e.,
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D=\{(\lambda_{i},z_{i}),f(\lambda_{i},z_{i}\cdot\mathds{1}_{b})\}_{i=1\text{\,with\,}1\leq z_{i}\leq b_{\text{max}}}^{N}$} (2)
Note that since $\mathds{1}_{b}$ acts as a constant scaling factor, and we will normalize $z_{i}$ w.r.t. $b_{\text{max}}$, we consider $\{(\lambda_{i},t_{i}),f(\lambda_{i},t_{i})\}_{i=1\text{\,with\,}0\leq t_{i}\leq 1}^{N}$ as dataset when discussing our surrogate.
@@
From a Bayesian perspective, we want to generate data from a distribution that captures our prior beliefs on the relationship between hyperparameters $\lambda$, training time $t$, and model performance $f(\lambda,t)$. While one could design such prior for a specific HPO scenario, our goal here is to construct a generic prior, resulting in an FT-PFN surrogate for general HPO. To this end, we leverage general beliefs, e.g., we expect learning curves to be noisy, but to exhibit an improving, convex, and converging trend; curves on the same task are expected to have similar start, saturation, and convergence points; and training runs using similar hyperparameter settings to produce similar learning curves.
@@
Following, Klein et al. (2017) and Kadra et al. (2023), we model the relationship between training time $t$ and $f(\lambda,t)$ by a parametric learning curve model $\hat{f}(t;\mathcal{E})$, whose parameters $\mathcal{E}$ are the output of a neural network $p(\lambda;\theta)$ taking the hyperparameter settings $\lambda$ as input (see Figure 2).
Figure 2: Diagram for the prior data model described in Section 4.1 that was used to generate data for meta-training FT-PFN. On the left, we have the randomly initialized neural network $p$ that models the relationship between a hyperparameter setting $\lambda$ and its learning curve (shown in pink), whose output parameterizes a curve model $\hat{f}$ that is a linear combination of $K$ (=2 in this illustration) basis functions (shown in red and blue) with added $\lambda$-specific Gaussian noise with variance $\sigma^{2}$.
@@
However, unlike these previous works, we do not train the parameters $\theta$ of this neural network. Instead, we randomly initialize the network, to represent a task-specific relationship between hyperparameters and their learning curves, which we then use to generate data for training FT-PFN. This can be viewed as generating samples from a Bayesian Neural Network (BNN) prior, meta-training FT-PFN to emulate LCNet-like BNN inference through in-context learning. As for $p$, we adopt the same architecture and initialization procedure for the BNN prior used by Müller et al. (2023) in the context of $B$-BO. The only difference here is that $p$ has $|\mathcal{E}|$ outputs, one for each parameter of our curve model $\hat{f}$, as opposed to a single output (model performance after a full training run). As for $\hat{f}$, following Domhan et al. (2015), Klein et al. (2017) and Adriaensen et al. (2023), we model learning curves as a linear combination of $K$ basis functions $f_{k}$, with additive Gaussian noise, i.e.,
$\begin{split}\hat{f}(t,\mathcal{E})\sim\mathcal{N}\left(\hat{f}_{comb}(t,\mathcal{E}),{\sigma^{2}}\right)\text{\quad with}\\ f_{comb}(t,\mathcal{E})=y_{0}+(y_{\infty}-y_{0})\cdot\sum_{k=1}^{K}w_{k}f_{k}(t,\Psi_{k})\\ \text{and \quad}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\lambda;\theta)=\mathcal{E}=\left(\sigma,y_{\infty},w_{1},\ldots,w_{K},\Psi_{1},\ldots,\Psi_{K}\right)}\end{split}$ (3)
@@
where $y_{0}$ is the model performance after 0 training steps, and $y_{\infty}$ that at convergence. [^1] Note that $y_{0}\notin\mathcal{E}$ as we assume it to be independent of $\lambda$. $w_{k}$ the weight of basis curve $f_{k}$, and $\Psi_{k}$ its the basis function specific parameters. Here, we adopt four different basis functions ($K=4$), each having four parameters, resulting in a total of 22 ($=|\mathcal{E}|$) parameters depending on $\lambda$ through pp. Our four basis functions subsume the power law model used by Kadra et al. (2023), all three basis functions used by Adriaensen et al. (2023), and 9 of the 11 basis functions originally proposed by Domhan et al. (2015). [^2] Our model excludes the two unbounded basis curves. Furthermore, unlike those considered in previous works, our basis functions can have a breaking point (Caballero et al., 2023) at which convergence stagnates or performance diverges, resulting in a more heterogeneous and realistic model.

### v2

@@ 1 Introduction
Hyperparameters are essential in deep learning (DL) to achieve strong model performance. However, due to the increasing complexity of these models, it is becoming more challenging to find promising hyperparameter settings, even with the help of hyperparameter optimization (HPO, Section 3.1) tools (Feurer & Hutter, 2019; Bischl et al., 2023). Traditional HPO techniques using Bayesian Optimization (BO) are unsuitable for modern DL because they treat the problem as a black box, making it computationally expensive, requiring a full model training for each evaluation.
Recent research in HPO has shifted towards multi-fidelity methods (Li et al., 2017, 2020a; Falkner et al., 2018; Klein et al., 2020; Li et al., 2020b; Awad et al., 2021), utilizing lower fidelity proxies (e.g., training for fewer steps, using less data, smaller models) and only evaluating the most promising hyperparameter settings at the full fidelity. While these methods have potential, they often use coarse-grained fidelity spaces and rely on the rank correlation of performances across fidelities. Moreover, they do not always fully utilize the anytime nature of algorithms such as checkpointing, continuation, and extrapolation. As a result, these methods struggle to allocate computational resources efficiently, leading to suboptimal performance in many scenarios.
The freeze-thaw BO method (Section 3.2), originally proposed by Swersky et al. (2014), is a promising approach to efficiently allocate computational resources by pausing and resuming the evaluation of different hyperparameter configurations. This method is an improvement over traditional grey-box methods as it dynamically manages resources, and recent implementations (Wistuba et al., 2022; Kadra et al., 2023) hold the state-of-the-art in the low-budget regime ($\sim$ 20 full function evaluations). However, these contemporary implementations have limitations. In particular, they rely on online learning to update the surrogate model at each step which can lead high computational overhead, instability, and complexity in managing additional hyper-hyperparameters. Moreover, they also suffer from strong assumptions about learning curves, which may not be applicable in all scenarios, resulting in overly confident incorrect predictions.
@@
In this work, we leverage prior-data fitted networks (Müller et al., 2022, PFNs) (Section 3.3), a Transformer-based meta-learning approach to Bayesian inference, to enhance freeze-thaw BO through in-context learning. Our PFN model (FT-PFN) infers the task-specific relationship between hyperparameter settings and their learning curves in a single forward pass, eliminating the need for online training during the search. Figure 1 compares learning curves extrapolation, including uncertainty, by our model (FT-PFN) and two baselines. Beyond demonstrating superior extrapolation quality, our model directly addresses key challenges in traditional HPO methods, lowering computational overhead and stabilizing the optimization process. The contributions of this paper are as follows:
  • We propose FT-PFN (Section 4.1), a new surrogate model for freeze-thaw BO, replacing online learning with in-context learning using PFNs. We train a single instance of FT-PFN model exclusively on synthetic data, generated from a curve prior designed to mimic realistic HPO learning curves.
@@
  • We combine FT-PFN with a novel acquisition function (MFPI-random, Section 4.2) and find that the resulting in-context freeze-thaw method (ifBO) yields a new state-of-the-art performance on three benchmark suites for HPO for deep learning (LCBench, Taskset, PD1).
The code for the surrogate PFN training and reproducing experiments from this paper, is available at: https://github.com/automl/ifBO.
@@ 2 Related Work
uses low-cost approximations of the objective function, for example, by evaluating only a few epochs of model training. A notable approach in this category is Hyperband (Li et al., 2017), which iteratively allocates a similar budget to various candidate hyperparameter settings and retains only the most promising ones, to be evaluated at a higher budget. Hyperband was extended for efficient parallelization (Li et al., 2020a), Bayesian optimization (Falkner et al., 2018) and evolutionary search (Awad et al., 2021) and to incorporate user priors (Mallik et al., 2023a). However, a key limitation of these multi-fidelity approaches is that they do not allow the continuation of previously discarded runs in the light of new evidence, resulting in a waste of computational resources. Additionally, their effectiveness heavily depends on a good manual choice of fidelity levels and a strong correlation between the ranks of hyperparameter settings at low and high fidelities (e.g., no crossings of learning curves).
@@
offers a promising solution to address the limitations of standard multi-fidelity and LCE-based HPO. Its ability to pause and resume optimization runs enables the efficient management of computational resources and ensures solid anytime performance. As a BO method, it incorporates a surrogate model to predict performance at higher fidelities and relies on the acquisition function to guide the search. The freeze-thaw concept was introduced by Swersky et al. (2014), who used a Gaussian process with exponential decay kernel for modeling learning curves and an entropy search acquisition function. Wistuba et al. (2022) recently proposed DyHPO, an improved version that uses a learned deep kernel combined with a multi-fidelity-based Expected Improvement (EI) acquisition. Most recently, Kadra et al. (2023) introduced DPL, another surrogate model that utilizes a deep ensemble of power laws and EI at the maximum budget as acquisition function. These Freeze-Thaw BO methods substantially improve upon standard multi-fidelity approaches, showing particularly strong performance in low-budget settings. Nevertheless, these methods share a common challenge: The necessity for online training of a surrogate model during the search, which can be computationally intensive and may cause optimization instabilities.
@@ 3.1 Hyperparameter Optimization (HPO)
Consider an iterative machine learning training pipeline with configurable hyperparameters $\lambda\in\Lambda$. E.g., when training a neural network using gradient descent we could configure the learning rate, weight decay, dropout rate, etc. Let $\operatorname*{f}(\lambda,b_{\lambda})$ be some measure of downstream performance (e.g., validation accuracy) of the model obtained, using hyperparameter settings $\lambda$, after $b_{\lambda}$ iterations of training, that we would like to maximize. HPO aims to find a hyperparameter setting producing the best model, i.e., $\lambda^{*}$: $(\lambda^{*},\cdot)\in\operatorname*{arg\!max}_{\lambda\in\Lambda,1\leq b\leq B}\operatorname*{f}(\lambda,b)$, within the limited total optimization budget of $B$ iterations. Note that in modern deep learning the budget available for HPO often does not allow us to execute more than a few full training runs. In this setting, the crux of HPO lies in allocating these limited training resources to the most promising hyperparameter settings, i.e., to find an allocation $\{b_{\lambda}\}_{\lambda\in\Lambda}$ with $b_{\lambda}\geq 0$ and $\sum_{\lambda\in\Lambda}b_{\lambda}\leq B$ that maximizes $\max_{\lambda\in\Lambda,\,1\leq b\leq b_{\lambda}}\operatorname*{f}(\lambda,b)$.
3.2 Freeze-Thaw Bayesian Optimization
@@
As discussed above, Swersky et al. (2014) proposed to address this challenge by following a fine-grained dynamic scheduling approach, allocating resources to configurations (and observing their performance) “one step at a time”. Here, one “step” corresponds to $\operatorname*{\mathds{1}_{b}}\geq 1$ iterations of model training (e.g., one epoch). Further, assume the maximum number of steps allocated to any single configuration $\lambda$ to be limited to $\operatorname*{b_{\text{max}}}$ (i.e., training runs are limited to $\operatorname*{b_{\text{max}}}\cdot\operatorname*{\mathds{1}_{b}}$ iterations $\approx$ training compute per configuration). Algorithm 1 shows the freeze-thaw Bayesian optimization framework that uses its history $H$, i.e., the various partial learning curves observed thus far in the current partial allocation, to fit a dynamic Bayesian surrogate model $\mathcal{M}{}$ that probabilistically extrapolates the partially seen performance of configuration $\lambda$ beyond $b_{\lambda}$ (Algorithm 1, line 6). Following the BO framework, it decides which of these to continue (or start if $b_{\lambda}=0$) using a dynamic acquisition policy $\mathcal{A}{}$ trading off exploration and exploitation (Algorithm 1, line 7). The crucial difference with traditional black box Bayesian Optimization lies in that resources in the freeze-thaw framework are allocated one step, rather than one full training run, at the time. Note, this framework assumes training to be preemptive, i.e., that we can stop (freeze) a model training, and continue (thaw) it later; this assumption is reasonable in modern DL where checkpointing is common practice. Also note, that black box BO can be recovered as a special case, by setting $\operatorname*{b_{\text{max}}}$, the maximum number of steps allocated to any single configuration, to one, and $\operatorname*{\mathds{1}_{b}}$ to the maximum budget available for any single training run.
For simplicity, in the remainder, we assume the budget step $\operatorname*{\mathds{1}_{b}}$ is set to 1, and the output of $\operatorname*{f}$ to be bounded in $[0,1]$.
@@
   $\operatorname*{f}$: measure of model performance to be maximized,
@@
   $\operatorname*{b_{\text{max}}}$: maximal steps for any configuration $\lambda\in\Lambda$,
@@
Output: $\lambda^{*}$$\in\Lambda$, obtaining the best observed performance
Procedure: HPO($\Lambda$, $\operatorname*{f}$, $\operatorname*{\mathds{1}_{b}}$, $\operatorname*{b_{\text{max}}}$, $B$):
1:  $b_{\lambda^{\prime}}\leftarrow 0$, $\forall\lambda^{\prime}\in\Lambda$
@@
4:
5:  while $|H|\cdot\mathds{1}_{b}<B$ do
6:   Train $\mathcal{M}{}$ on $H$ FT-PFN requires no model fitting
7:   $\lambda\xleftarrow{}$ $\mathcal{A}{}$($\Lambda,\mathcal{M}{},H,\operatorname*{b_{\text{max}}}$) select $\lambda$ to thaw
8:   $b_{\lambda}\leftarrow b_{\lambda}+\mathds{1}_{b}$thaw $\lambda$ for one step
9:   $y\leftarrow{}\operatorname*{f}(\lambda,b_{\lambda})$evaluate $\operatorname*{f}$
10:   $H\leftarrow H\cup\{(\lambda,b_{\lambda},y)\}$
11:  end while
11:
12:  return $\lambda^{*}$: $(\lambda^{*},\cdot)\in\operatorname*{arg\!max}_{\lambda\in\Lambda,\,1\leq b\leq b_{\lambda}}\operatorname*{f}(\lambda,b)$
@@ 3.3 Prior-data Fitted Networks (PFNs)
As briefly discussed in Section LABEL:sec:icl, PFNs (Müller et al., 2022) are neural networks $q_{\theta}$ that are trained to do Bayesian prediction for supervised learning in a single forward pass. More specifically, let $D=D_{\text{train}}\cup\{(x_{\text{test}},y_{\text{test}})\}$ be a dataset used for training; the PFN’s parameters $\theta$ are optimized to take $D_{\text{train}}$ and $x_{\text{test}}$ as inputs and make predictions that approximate the posterior predictive distribution (PPD) of the output label $y_{\text{test}}$:
$q_{\theta}(x_{\text{test}},D_{\text{train}})\approx\mathbb{P}(y_{\text{test}}\,|\,x_{\text{test}},D_{\text{train}}),$
in expectation over datasets $D$ sampled from a prior $p(\mathcal{D})$ over datasets. At test time, the PFN does not update its parameters given a training dataset $D_{\text{train}}$, but rather takes $D_{\text{train}}$ as a contextual input, predicting the labels of unseen examples through in-context learning. The PFN is pretrained once for a specific prior $p(\mathcal{D})$ and used in downstream Bayesian prediction tasks without further fine-tuning. More specifically, it is trained to minimize the cross-entropy for predicting the hold-out example's label $y_{\text{test}}$, given $x_{\text{test}}$ and $D_{\text{train}}$:
$\begin{split}\ell_{\theta}=\mathbb{E}[\shortminus\text{log}\ q_{\theta}(y_{\text{test}}|x_{\text{test}},D_{\text{train}})]\\ \text{with \quad}\{(x_{\text{test}},y_{\text{test}})\}\cup D_{\text{train}}\sim p(\mathcal{D})\end{split}$ (1)
@@ 4 In-Context Freeze-Thaw BO (ifBO)
In this section, we describe ifBO, the in-context learning variant of the freeze-thaw framework that we propose as an alternative for the existing online learning implementations (Wistuba et al., 2022; Kadra et al., 2023). As can be seen in Algorithm 1 (line 6), the critical difference lies in the fact that we do not need to refit our surrogate model after every allocation step. By skipping the online refitting stage, we reduce computational overhead, code complexity, and hyper-hyperparameters. Note that within the freeze-thaw BO framework described in Section 3.2, our method is fully characterized by our choice of surrogate model (Section 4.1) and dynamic acquisition policy (Section 4.2).
@@ 4.1 Dynamic Surrogate Model (FT-PFN)
We propose FT-PFN a prior-data fitted network (Müller et al., 2022, PFNs) trained to be used as an in-context dynamic surrogate model in the freeze-thaw framework. As described in Section 3.3, PFNs represent a general meta-learned approach to Bayesian prediction, characterized by the data used for meta-training. Following previous works using PFNs (Müller et al., 2022; Hollmann et al., 2023; Müller et al., 2023; Adriaensen et al., 2023; Dooley et al., 2023), we train the PFN only on synthetically generated data, allowing us to generate virtually unlimited data and giving us full control over any biases therein. We aim to train FT-PFN on artificial data ($D$) that resembles the real performance data we observe in the context of freeze-thaw Bayesian optimization, i.e., a collection of learning curves of training runs for the same task, but using different hyperparameter settings, i.e.,
$\bigcup_{\lambda\in\Lambda}\biggl\{\bigl((\lambda,1),\operatorname*{f}(\lambda,1)\bigl),\ldots,\bigl((\lambda,\operatorname*{b_{\text{max}}}),\operatorname*{f}(\lambda,\operatorname*{b_{\text{max}}})\bigl)\biggl\}$
@@
From a Bayesian perspective, we want to generate data from a prior data model that captures our beliefs on the relationship between hyperparameters $\lambda$, training budget $b$, and model performance $\operatorname*{f}(\lambda,b)$. While one could design such prior for a specific HPO scenario, our goal here is to construct a generic prior, resulting in an FT-PFN surrogate for general HPO. To this end, we leverage general beliefs, e.g., we expect learning curves to be noisy, but to exhibit an improving, convex, and converging trend; curves on the same task are expected to have similar start, saturation, and convergence points; and training runs using similar hyperparameter settings to produce similar learning curves.
Figure 2: Diagram for the prior data model described in Section 4.1 that was used to generate data for meta-training FT-PFN. On the left, we have the randomly initialized neural network $\operatorname*{\pi_{\text{config}}}$ that models the relationship between a hyperparameter setting $\lambda$ and its learning curve (shown in pink), whose output parameterizes a curve model $\operatorname*{\pi_{\text{curve}}}$ that is a linear combination of $K$ (=2 in this illustration) basis functions (shown in red and blue) with added $\lambda$-specific Gaussian noise with variance $\sigma^{2}$.
@@
Following Klein et al. (2017) and Kadra et al. (2023), we model the performance curve of a hyperparameter $\lambda$ using a parametric curve model $\operatorname*{\pi}_{\text{curve}}$, whose parameters are sampled from an another prior model $\operatorname*{\pi}_{\text{config}}$ taking the hyperparameter $\lambda$ as input (see Figure 2). Following Domhan et al. (2015) and Adriaensen et al. (2023), we define $\operatorname*{\pi_{\text{curve}}}$ as a weighted combination of $K$ basis functions $f_{k}$, with additive Gaussian noise. Thus, the parameters of $\operatorname*{\pi_{\text{curve}}}$ include $\mathcal{E}$ (the set of parameters and weight of all basis functions), along with $\sigma^{2}$ (noise). As for $\operatorname*{\pi_{\text{config}}}$, we adopt a neural network with weights $\theta$. Unlike previous works, we do not train the weights $\theta$ of this neural network. Instead, we randomly initialize the network, to represent a task-specific relationship between hyperparameters and their learning curves, which we then use to generate data for training FT-PFN. This can be viewed as generating samples from a Bayesian Neural Network (BNN) prior, meta-training FT-PFN to emulate LCNet-like (Klein et al., 2017) BNN inference through in-context learning.
Formally, we define the performance of a hyperparameter $\lambda$ at a training time $t$ as follows:
$\operatorname*{\pi_{\text{curve}}}(\lambda,t)\sim\mathcal{N}\left(f_{\text{comb}}(t;\mathcal{E}),\sigma^{2}\right)$ (2)
$\text{with}\quad f_{\text{comb}}(t;\mathcal{E})=y_{0}+(y_{\infty}-y_{0})\cdot\sum_{k=1}^{K}w_{k}f_{k}(t;\Psi_{k})$
$\text{and}\quad(\sigma^{2},\underbrace{\left(y_{\infty},w_{1},\ldots,w_{K},\Psi_{1},\ldots,\Psi_{K}\right)}_{\mathcal{E}})\sim\operatorname*{\pi_{\text{config}}}(\lambda;\theta),$
@@
where $y_{0}$ is the initial model performance [^1] Note that $y_{0}\notin\mathcal{E}$ as we assume it to be independent of $\lambda$., and $y_{\infty}$ that at convergence. $w_{k}$ is the weight of basis curve $f_{k}$, and $\Psi_{k}$ its basis function specific parameters. In this work, we adopt four different basis functions ($K=4$), each having four parameters, resulting in a total of 22 ($|\mathcal{E}|+1$) parameters depending on $\lambda$ through $\operatorname*{\pi_{\text{config}}}$. Our four basis functions subsume the power law model used by Kadra et al. (2023), all three basis functions used by Adriaensen et al. (2023), and 9 of the 11 basis functions originally proposed by Domhan et al. (2015). [^2] Our model excludes the two unbounded basis curves. Furthermore, unlike those considered in previous works, our basis functions can have a breaking point (Caballero et al., 2023) at which convergence stagnates or performance diverges, resulting in a more heterogeneous and realistic model.